Final Concept Quiz Section Full Name: Higher Dimensions 1) Which combination of options is most accurate (A) Acceleration always means an objects speed is increasing or decreasing (B) When the net force is perpendicular to momentum, the object will turn. (C) A projectiles time in the air is fully determined by its initial horizontal velocity (D) Change in momentum is directed in the same direction as the net force. (E) None of the above Force and Energy 2) Which combination of options is most accurate (A) Equilibrium corresponds to when the potential energy of a system is zero. (B) The total reversible force on a system is zero when its potential energy is at a minimum. (C) Work is only done on a system if the systems kinetic energy increases. (D) Work only occurs if the forces go through a displacement. (E) None of the above orce and Momentum
Answer #1

1.

A.Accelaration is non zero is uniform circular motion ,but there is no change in speed

B.Object have to turn because force will create a velocity component perpendicular to the initial motion and therefore it have to turn

C.No,T=2usin(a)/g ,clearly usin(a) is vertical component of velocity

D.Yes,F=dp/dt

So b,d are correct

2.

A.Not necessarily,but F=-dU/dx =0

B.Correct but not accurate because F=- dU/dx =0 for Potential Energy minimum maximum or U-x curve has a point of inflection

C.Nope,work done on system can increase Potential Energy too

D.Yes,cause W=

\int F.dr

So D is correct